OpenStudy (anonymous):

16(x^2+y^2)^2=400xy^2 ; (4,2) write an equation for the tangent line

OpenStudy (anonymous):

I somewhat get it.. so do i distribute the 16 to x^2 and y^2 and get the dev. of that?

OpenStudy (zzr0ck3r):

you cant distribute until you expand the binomial there are two ways of going about it, you can expand the binomial then distribute the 16 or you can use the chain rule

OpenStudy (zzr0ck3r):

on the RHS you will need to use the product rule and chain rule and remember this is implicit differentiation.

OpenStudy (anonymous):

2(x^2 +y^2) * 2x + 2y *y'?

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

well 2(x^2 +y^2) * (2x + 2y *y')

OpenStudy (anonymous):

And then I move everything besides the y' to the right?

OpenStudy (zzr0ck3r):

you need to implicite differentiate the right hand side

OpenStudy (zzr0ck3r):

implicit*

OpenStudy (anonymous):

y'= 2(x^2+y^2)/2x+2y like so?

OpenStudy (zzr0ck3r):

you need to implicit differentiate the right hand side

OpenStudy (zzr0ck3r):

\(400xy^2\implies400(y^2+x2yy')\)

OpenStudy (zzr0ck3r):

now we have \(2(x^2 +y^2) * (2x + 2y *y')=400(y^2+x2yy')\)

OpenStudy (zzr0ck3r):

solve for y'

OpenStudy (zzr0ck3r):

its not going to be fun.....

OpenStudy (zzr0ck3r):

you with me still?

OpenStudy (anonymous):

yup im trying to solve for y' now!

OpenStudy (zzr0ck3r):

show me what you get, ill do it on paper:)

OpenStudy (anonymous):

in this kind of problem you don't have to move x's and y's around. just plug in the numbers you have for x and y, then solve for y'

OpenStudy (anonymous):

i'm not the best at this yet, so it might take some time! take you for helping me [:

OpenStudy (zzr0ck3r):

np i got \(y'=\frac{400y^2-4x^3-4xy^2}{-800xy+4x^2y+4y^3}\)

OpenStudy (anonymous):

@zzr0ck3r can I do that? instead of adding, subtracting, and dividing then plugging in the numbers can I just plug in the numbers for x and y and then work with only y' as a variable?

OpenStudy (zzr0ck3r):

but its late. after this you will plug in your point for (x,y) and that will give you the slope of the line y=mx+b then use the point again to find b and we are done \(\huge yay!!!!\)

OpenStudy (zzr0ck3r):

sure, if its the one time thing

OpenStudy (zzr0ck3r):

but sometimes it comes in handy to have the equation

OpenStudy (anonymous):

isn't point-slope form easier than y=mx+b to get the equation?

OpenStudy (anonymous):

thank you so much for the help! :D

OpenStudy (anonymous):

if you know how to use it?

OpenStudy (zzr0ck3r):

i.e if you have to evaluate more than one point, then find an explicit equation.

OpenStudy (zzr0ck3r):

sure if its easier to you, there are both the same to me.. I always use that one because every knows it.

OpenStudy (anonymous):

what i got for that was -.0979487 ... sorry i am so slow lol

OpenStudy (anonymous):

y=-.0979487x+-1.91795?

OpenStudy (ranga):

16(x^2+y^2)^2=400xy^2 Differentiate both sides implicitly: 32(x^2 + y^2)(2x + 2yy') = 400(2xyy' + y^2) To evaluate y' at (4,2), put x = 4 and y =2 right here and then find y' 32(16 + 4)(8 + 4y') = 400(16y' + 4) 32*20*4*(2 + y') = 400*4*(4y' + 1) 32(2+y') = 20(4y'+1) 8(2+y') = 5(4y'+1) 16+8y' = 20y' + 5 11 = 12y' y' = 11/12 at the point (4,2) y = mx + b y = 11/12x + b x = 4, y = 2 2 = 11/12 * 4 + b b = 2 - 11/3 = (6-11)/3 = -5/3 y = 11x/12 - 5/3 is the equation of the tangent line.